Find the mean of the probability distribution of a random variable X which can take 

only the values 2, 4, 5, and 9, given that P(2)=9/20, P(5)=1/20, P(5)=1/5, and 

P(9)=3/10


1
Expert's answer
2022-03-22T18:59:31-0400

"EX=2\cdot \frac{9}{20}+4\cdot \frac{1}{20}+5\cdot \frac{1}{5}+9\cdot \frac{3}{10}=4.8"